OpenStudy (anonymous):

What is the exact area of a circle with a diameter of 37 inches?

OpenStudy (anonymous):

Remember this: \[A=\pi r^2\] r is half the diameter

OpenStudy (richyw):

\(A=\pi r^2=\pi \frac{d^2}{4}\)

OpenStudy (asnaseer):

You have already posted this question earlier here: http://openstudy.com/updates/500ad2fae4b0549a892f282f where @maheshmeghwal9 helped you out.